There are a total of 100 men and women on an airplane.There are 12 more women than men. What is the number of women on the airplane?
Answer by ikleyn(52775)

.

men + women    = 100


men + (men+12) = 100


2*men + 12     = 100


2*men          = 100-12


2*men          = 88


men            = 88/2 = 44.    ANSWER

Solved.